1. Determine the general solution of the equation: 4sin2 x - 3 = 0
1
Expert's answer
2021-07-04T15:48:36-0400

"4\sin^2x-3=0\\
4\sin^2x=3\\
\sin^2x=\frac{3}{4}\\
\sin{x}=\pm{\sqrt{\frac{3}{4}}}\\
\sin{x}=\pm\frac{\sqrt{3}}{2}\\
x=\sin^{-1}{(\frac{\sqrt{3}}{2})}\\
x=60\degree, 120\degree\\
x=\sin^{-1}{(-\frac{\sqrt{3}}{2})}\\
x=300\degree, 240\degree\\"

General solution is

"x=360\degree(n)+60\degree, 360\degree(n) +120\degree, 360\degree(n)+240\degree, 360\degree(n)+300\degree"